PLEASE ANSWER ALL THE QUESTIONS

1. List the parts of a light microscope and include the specific use of each part. Give one (3 points)

2. The MW of Trizma base is 110. How much of this chemical you need to make 200 ml of 0.2 M solution? Show calculations and briefly describe as to how you will make this solution and adjust pH to 8. (2 points)

3. Given the molecular weight of Glucose 180, answer the following questions. Show calculations. (3 points)

a. How much glucose you need to make 250 ml of 2 M solution?

b. How will you make 100 ml 0.5 M solution from the above 2 M stock solution?

4. How will you do serial dilution to make 10 ml each of 50 mM, 5 mM, 0.5 mM solutions from the above 0.5 M solution?
Answer #1

1. A light microscope is an instrument used in Biology lab to visualise microscopic organisms which cannot be seen with the help of a naked eye.

The parts of a microscope are -

Eyepiece/occular lens - the lens of microscope through which the viewers views into the specimen with the help of his eyes

Objective lens - the lens of microscope which is is closest to the specimen being viewed

The purpose of both the lenses is to magnifies the specimen being viewed

Diopter adjustment - changes the focus of eyepieces per the eyesight of the viewer

Head - the path connecting eyepiece with objective lens

Frame - connects body tube to the base of microscope

Stage control - control the horizontal and vertical movement of the stage

Coarse adjustment knobs - adjust the focus of the specimen

Fine adjustment knobs - increases the details of the specimen

Light switch - used to switch on and off the light source

Base - the path at which hole assembly of microscope stands

Brightness adjustment - increases or decreases the intensity of light entering the specimen

Condensor - collect and focuses light from light source into the specimen

Mechanical stage - stage on which the the slide or specimen is kept

Nosepiece - rotating part which holds all the objective lenses together

For questions, below use formula,

Molarity (M) = given mass (m)/molecular mass (MW) Volume in L (V)

2.

MW = 110

M = 0.2

V = 200/1000 = 0.2 L

0.2 = m/110(0.2)

m = 44

Steps to prepare the solution -

A. Weigh 44 g of trizma using a weighing balance.

B. Dissolve it in around 50 mL of water after proper shaking.

C. Now make the final volume to 200 mL.

D. Using, a pH meter, adjust the pH of th solution. If the pH of the solution is more than 8 then add acid or if it is less then 8 then add a base.

3.

A. 2 = m/180(0.25)

m = 2 × 180 × 0.25 = 90 g

90 g glucose is required to make 2 M 250 mL solution.

B. To make solution from stock solution, use formula,

Volume of stock required × molarity of stock = volume of solution to be made × molarity of solution required

Volume of stock required × 2 = 100 × 0.5

Volume of stock required = 25 mL

To make 100 mL, 0.5 M solution, add 25 mL 2M solution in 75 mL of distilled water.

4.

Always remember to take the volume in litres only. Convert mL to L. Or else you will get false results.

Please rate.